From: william@physics.utoronto.ca (William Trischuk)
Date: Sat, 28 Feb 2004 05:58:15 -0500 (EST)



UEC Minutes -- February 21, 2004
---------------------------------

Present: Bloom, Garcia, Griffiths(GSA), Gottschalk, Hagopian, Lobo(GSA),
Rolli, Sheldon(phone), Tanaka, Tschirhart, Trischuk, White,
Zimmerman, apologies: Groer, Messier.


User's Meeting Planning (White)

- The Users Meeting is scheduled for June 2nd and 3rd, 2004. Our
sponsors for the meeting (the URA) are pushing us to stick much
closer to the budget for the meeting this year. We are considering
a number cut-backs to stay on budget including: not paying for
drinks at the dinner, reducing the scope of the video/audio
archiving that is done for the meeting and perhaps asking users to
pay $5/person for the dinner. We hope to converge on this in the
coming weeks to find a workable budget.

- Mike Turner, Irving Lerch have agreed to make presentations at
the meeting. We are now trying to arrange for a visit from the
Director General of CERN (Aymar) and the KEK Director
(Totsuka). Still working on getting representatives from the
funding agencies as we have managed to pick a date that conflicts
with a SLAC DOE review. Should be no problem filling the agenda.

- Thesis competition and Postdoc competitions are on track but so
far there have been few submissions -- need to work on spreading
the word.


Status of the Laboratory (Mike Witherell)

- Preparing for a DOE review of the accelerator complex during the
week of February 23rd. 17 reviewers formed into six subcommittees
will visit for three days. There are good prospects to double the
luminosity delivered up to FY03 during FY04. The "plan" is the
design that will deliver 0.31 fb^-1 in FY04. We are now two weeks
ahead of the plan. The work in the Fall '03 shutdown has taken us
from 27 pb^-1 per month to somewhere beyond 40 pb^-1 per
month. All records have be broken in calendar year 2004. The only
cloud on the horizon is that the increased anti-proton stacking
hasn't been achieved. This will be addressed in a one-week
shutdown during March. The slides from the upcoming accelerator
review will be available on the web at:
http://www-bdnew.fnal.gov/doereview04/index.htm

- Collider programme is delivering increased luminosity. CDF has
seen aging in their wire-chamber. They plan is to increase gas
flow and perhaps change the gas mixture to abate the aging, but in
a one-week shutdown in March they will replace a wire-plane and
examine the old plane to see if the cause of the aging can be
understood. This will be done in conjunction with the anti-proton
source work as well as further alignment of low beta
quads. Mini-Boone also received its best month of protons in
January, but still work is still needed to increase the rate of
putting protons on target including the positioning of Booster
collimators to reduce the losses. These losses/collimators are now
the focus of the Accelerator divisions 'rapid response' team. The
switch-yard and fixed target extraction have been commissioned
after working through alignment issues. Fast spill has been
achieved and now they are working on slow-spill for testbeam work.

- The CD-0 for BTeV was signed in mid-February. This is the first
step on the path to completion at CD-4 (CD-3 is approval for
construction spending). CD-1 review planned for the end of
April. Director's review of the C0 interaction region was
conducted in February -- shows that there are good prospects to
have a beta-star at C0 comparable to that of the current collider
experiments. Operations review of the lab is being prepared to
make sure that the expected funding will allow the existing
programme to continue while the work on BTeV construction gets
underway.

- Q: What is the current schedule for Linear Collider technology
decision? A: The Director didn't attend Paris LC International
steering meeting in early February but there the 12-person
technical decision panel (chaired by Barrish) had its first
meeting. They have access to the cold/warm cost comparison that
has been done in the US. This comparison will be made public as
soon as the relevant lab directors have been consulted --
hopefully by April. The 12-person committee is charged with making
its decision "this year". There is still some hope it may be as
early as Fall 2004. Q: HEPAP expressed some concern that LC was
meeting. A: Yes, all the things in the first group were on their
way to CD-0 (like BTeV). The LC was not as advanced and so the
best it could make was make the second group, where it appears as
the highest priority facility. The LC also does not appear in the
DOE budget summaries. At the HEPAP meeting Orbach said that the
inclusion of the Linear Collider in the facilities plan was an
important statement of its scientific priority. He implied that
any project that large would need to get special support that
highest level of government for a construction start.

Q: Is the FNAL long range planning report due out soon? A: The
committee is converging. The forum for the report's release is
being discussed. The URA board of overseers met recently and heard
a draft version. The PAC will receive an updated draft version
soon. Should be available in plenty of time for the User's
Meeting in June.

Q: Is the all experimenters meeting the right forum to bring up
concerns over proton intensity being delivered to miniBooNE?
A: Yes, but he is aware that nothing has been getting as much
attention as the collider program, especially in the period
leading up to the DOE review. Assigning the Rapid Response Team in
the accelerator division to the problem is an important step. The
accelerator division will also be developing a set of measures of
Booster performance to track progress, much as is done for collider
luminosity.


Overview of Beam Instrumentation at Fermilab (Bob Webber)

- The slides from his talk are available at:

http://beamdocs.fnal.gov/DocDB/0010/001033/001/UECtalk_02_2004.pdf

- There are a wide range of devices being used to monitor beams in
machines and transfer lines at the lab. Apply the information to know
how well the machines are working, use it for real time feedback
to make the machine work better and to try to understand how the
magnets are working and the machine tunes. Things that are measured
include: current (intensity), position of the beams, size of the beams
and beam losses.

- Customers include Tevatron/RunII programme which absorb the
lion's share of the operational support. Considerable effort going
into upgrades of the PBar transfer line monitoring and Tevatron
abort gap loss monitors. His group is also working on upgrades to
the beam loss monitors that will contribute to a more robust Tevatron
abort system. Have been re-commissioning old systems in testbeam
extraction to the Meson Hall. NUMI is getting a set of new
monitors as they approach first beam later this year.

- Provided a list of priorities from his perspective includes
identifying the objectives and getting devices ready for NuMI's
first beam. Tevatron beam position monitor (BPM) data acquisition
system is being upgraded with help from the computing
division. PBar transfer lines, BPMs and Tevatron beam loss
monitors are also high on his list. The Particle Physics Division
is contributing effort in numerous areas including the Tevatron
loss monitors.

- Showed some results from prototype new Tevatron beam position
monitor electronics with dramatically improved resolutions and
significantly reduced systematics. Resolutions better than 10
microns have been demonstrated (current system has a least count
of 150 microns). Showed results of turn-by-turn studies of beams,
that have been given explicit orbit kicks.


Implementation of Proton Source Recommendations (Eric Prebys)

The slides from his talk are available at:

http://home.fnal.gov/~prebys/talks/uec_20040221.ppt

- Following up on the Finley report that highlighted the "proton
crunch" that is coming when NuMI comes online. Most of the proton
source (Cockroft Walton, original Linac, "new" Linac and Booster)
is basically original lab equipment dating back 30 years ("new"
Linac is 10 years old). The main injector is a key component for
NuMI -- relatively new. Booster can supply 5E12 protons at
7.5Hz. Would have to go to 10Hz to supply both BooNE and
NuMI. Loading the MI takes about 1.4s setup + 1/15s from one
Booster cycle. The Booster operation/rep rate is currently limited
by beam losses and the resulting activation of machine
components. If components become too activated they can't be
maintained. The number of protons that can be accelerated is
currently limited by a limit of 400W of overall power loss around
the ring.

- Discussed the widely distributed "proton demand" plot and how
they are working on reducing losses to get closer to the miniBooNE
request for protons. FY03 projects included improvements to the
Linac system (including cooling) and lots of work on the Booster
to mitigate losses (collimators, dogleg in the extraction region, etc.).

- Near time priorities: optimising lattice, further tuning of
collimator positions and increasing the aperture in key locations.
New pieces of the vacuum system were almost ready for the past
shutdown. They will go in soon. Further work by the Accelerator
Division rapid response team and alignment task force is expected
to help. Trying to satisfy conflicting demands of improved
alignment of machine elements and optimising the aperture. Learning
how to integrate surveys with known displacements that have been
made to Booster machine elements.

- For the next year want to further improve the Linac
instrumentation but the bulk of the work will be on the
Booster. Will also begin to focus on 'multi-batch' filling of Main
Injector (MI). Need to synchronise the Booster and MI timing to allow
multiple transfers per MI cycle. This kind of timing
synchronisation is also necessary for the slip-stacking of anti-protons.

- Developing a plan to maintain the current proton source through
FY08 under the assumption that a new proton source will be built
in the out-years. Has been given some budgetary guidance that
consists of spending $18M over the next 5 years. This is not
sufficient to consider replacing Linac front end or changing the
MI ramping time. Could rebuild the Booster RF and nothing
else. Could also consider adding two additional RF cavities. Hope
to have detailed plan by the summer of this year.

- Reasonable expectations: 1E13 protons per pulse for pBar
production; 5E20 protons to miniBooNE before NuMI turns on. Then
should be able to achieve 2 to 2.5E20p/yr for NuMI. If things go
well might be able to satisfy pBar, NuMI and still supply enough
protons to make it worthwhile continuing to run
miniBooNE. Supplying more than 4E20 p/yr seems unrealistic. Off
Axis neutrinos want at least twice this. This would require an
investment in the current system on the order of $100M, which
would directly compete for resources with a new proton driver
-- one of the projects being considered by the lab's long range
planning committee.


GSA report (Lobo)

- Shared a list of concerns of the GSA membership with the
UEC. These include: rooms on site for students doing owl shifts --
would require some arrangement with the housing office in the
village and ask the experiment spokespeople; investigating further
extensions of the taxi service off-site and after 4pm (might want
to consider inviting a PACE bus on-site if there is sufficient
interest); also concerned about the cafeteria's use of disposable
dishes and cutlery (and the lack of recycling on-site in general).
Finally there are still major concerns about visa issues and the
process of fingerprinting visitors who need visas to enter the US.

- Are in the process of planning the GSA sponsored "New
Perspectives" meeting that will follow the users meeting in
June. They are currently looking for speakers.


Planning for the UEC/SLUO DC trip (Zimmerman)

- 13 people from the UEC (and GSA) and 11 people from SLUO (the SLAC
Users Organisation) will visit Washington from March 24/25/26. The meeting
will start with visits to members of congress and the senate. We are making
every effort to visit members of all the relevant/important sub-committees.
An initial pass of assignments to make appointments, based on where people
live is now final, but a second pass is being prepared as there seems to
be time to visit with more members.


Next meeting March 20, 2004